martin and anna buy books at a sale. martin buys 3 hardcover books and 4 paperback books for $6.50. Anna buys 2 hardcover books and 6 paperback books for $6.00. What is the cost, in dollars, of each hardcover book?
Answer by checkley75(3666)

3h+4p=6.50
2h+6p=6.00 noe solve for h
2h=-6p+6.00
h=-6p/2+6.00/2
h=-3p+3.00 now substitute (-3p+3.00) for h in the first equation thus:
3(-3p+3.00)+4p=6.50
-9p+9.00+4p=6.50
-9p+4p=6.50-9.00
-5p=-2.50
p=-2.50/-5
p=.50 is the price for the paper back books.
now solve for h
3h+4*.50=6.50
3h+2.00=6.50
3h=6.50-2.00
3h=4.50
h=4.50/3
h=1.50 cost of the hard cover books.
proof
2*1.50+6*.50=6.00
3.00+3.00=6.00
6.00=6.00